OpenStudy (gabylovesyou):

A greenhouse that specializes in growing bell peppers is divided into sections. The number of plants in each section depends on the number of sprinklers in that section. In a section with x sprinklers, there are 3x(x + 3) red bell pepper plants and (x + 4)2 yellow bell pepper plants. Which simplified expression represents the total number of red and yellow bell pepper plants in a section with x sprinklers? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'm assuming (x + 4)2 is really (x + 4)^2, if so, then Add up 3x(x + 3) and (x + 4)^2

jimthompson5910 (jim_thompson5910):

3x(x + 3) + (x + 4)^2 3x(x + 3) + (x + 4)(x + 4) What's the next step?

OpenStudy (gabylovesyou):

yes it is :)

jimthompson5910 (jim_thompson5910):

how would you expand out (x + 4)(x + 4)

OpenStudy (gabylovesyou):

Foil ?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (gabylovesyou):

x * x = x^2 x *4 = 4x 4 * x = 4x 4 * 4 = 16

jimthompson5910 (jim_thompson5910):

add all that up

OpenStudy (gabylovesyou):

x^2 + 4x + 4x + 16 = x^2 + 8x + 16

jimthompson5910 (jim_thompson5910):

3x(x + 3) + (x + 4)(x + 4) 3x(x + 3) + x^2 + 8x + 16 what's next

OpenStudy (gabylovesyou):

ummmm

OpenStudy (gabylovesyou):

3x^2 is 3x(x + 3) right ?

jimthompson5910 (jim_thompson5910):

3x(x + 3) 3x*x + 3x*3 what's after this

OpenStudy (gabylovesyou):

3x^2 + 9x ?

jimthompson5910 (jim_thompson5910):

3x(x + 3) + x^2 + 8x + 16 3x^2 + 9x + x^2 + 8x + 16 then what?

OpenStudy (gabylovesyou):

combine like terms.... 4x^4 would be 3x^2 and x^2 right ?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

what about 9x and 8x

OpenStudy (gabylovesyou):

17x

jimthompson5910 (jim_thompson5910):

put that all together to get ???

OpenStudy (gabylovesyou):

wait 4x^4 or 4x^2 ?

jimthompson5910 (jim_thompson5910):

the x^2 terms stay the same

jimthompson5910 (jim_thompson5910):

so 3x^2 + x^2 = 4x^2

OpenStudy (gabylovesyou):

the answer is B :)

OpenStudy (gabylovesyou):

thank you!

jimthompson5910 (jim_thompson5910):

well I don't see the answer choices, so tell me what you got for the final answer

OpenStudy (gabylovesyou):

4x2 + 17x + 16

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

sry had to step away for a bit, but just wanted to check your final answer

OpenStudy (gabylovesyou):

i tagged u in my other question....they r soooo confusing can u explain better ? lol